# Tuning Event Camera Biases Heuristic for Object Detection Applications in Staring Scenarios

### v1

@@ 5 Proof of Theorem 1
$\underset{y\to c}{\lim}g(x,y)$ $=0$
$\underset{y\to d}{\lim}g(x,y)$ =$\infty$= ∞
and the monotony of $g$ with relation to $y$, by the intermediate value theorem, it follows that for every $a\leq x<b$ there exists exactly one value $y(x)$ such that
$g(x,y(x))=C.$ (5.1)
Now, we want to show that the function $f(x,y(x))$ is monotonically increasing as function of $x$, even though $y$ is not fixed.
Using the formula for the derivative of the implicit function $y(x)$ given by Equation 5.1, one has
$\frac{\partial y}{\partial x}=-\frac{\partial g}{\partial x}/\frac{\partial g}{\partial y}.$
Hence, by the chain rule and the assumptions of the theorem on the derivatives of $f$ and $g$, one has
$\frac{\partial f(x,y(x))}{\partial x}$ $=\frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}\cdot\frac{\partial y}{\partial x}$
$=\frac{\partial f}{\partial x}-\frac{\partial f}{\partial y}\cdot\left(\frac{\partial g}{\partial x}/\frac{\partial g}{\partial y}\right)\geq\frac{\partial f}{\partial x}-\frac{\partial g}{\partial x}>0$
yielding that $f(x,y(x))$ is monotonically increasing as required. Now, notice that
$\underset{x\to a}{\lim}f(x,y(x))$ $\leq\underset{x\to a}{\lim}f(x,d)=0$
$\underset{x\to b}{\lim}f(x,y(x))$ $\geq\underset{x\to b}{\lim}f(x,c)=\infty$
and therefore
$\underset{x\to a}{\lim}f(x,y(x))$ $=0$
$\underset{x\to b}{\lim}f(x,y(x))$ =$\infty$= ∞
Using the intermediate value theorem once again, for the monotonically increasing function $f(x,y(x))$, it follows that there exists exactly one $x$ such that
$f(x,y(x))=C.$
This gives us the first part of Theorem 1. For the second part, let us write the equations
@@
11 11 $=\frac{\partial f}{\partial C}=\frac{\partial f}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial f}{\partial x}\cdot\frac{\partial x}{\partial C}$
11 11 $=\frac{\partial g}{\partial C}=\frac{\partial g}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial g}{\partial x}\cdot\frac{\partial x}{\partial C}$
Hence
$\left(\begin{array}[]{cc}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}\\ \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}\end{array}\right)\cdot\left(\begin{array}[]{c}\frac{\partial x}{\partial C}\\ \frac{\partial y}{\partial C}\end{array}\right)=\left(\begin{array}[]{c}1\\ 1\end{array}\right)$
@@ 6 Determining filters biases - demonstration
Then, we sampled the plane spanned by $L=$ bias_fo and $H=$ bias_hpf in steps of 10 units. For each sample we used an automatic feedback control algorithm that balanced the positive and negative event rate to be close to 50k events/sec and 100k events/sec in total. We remark that due to Theorem 1, the convergence of the algorithm to the desired background event rate was guaranteed. Conceptually, the algorithm performs a similar idea as the AGC feature in frame based cameras.
@@ 7 Conclusions
Solving the biases multivariable problem is a crucial stage one needs to pass to exploit the potential of event camera for performing any task. Hence, developing systematic heuristics and approaches to tackle this problem is an important challenge of the research community.
In this paper we presented a systematic heuristic to characterize the biases of an event camera for detection of small objects tasks in staring scenarios. We translated the biases behavior in this context into mathematical terms and showed that under suitable and reasonable assumptions, for any value of the filters biases, there exists exactly one pair of sensitivity biases that provide a predefined positive and negative event rate.
Having this understanding in mind, we demonstrated, in a certain scenario, under appropriate constraints on the event rate, how the sensitivity biases change, and how the signal of an incandescent lamp powered by the electrical grid is being influenced as well. This demonstraion gave us a better understanding on what can be done with event camera when one wants to detect signals with similar behavior.

### v2

@@ 5 Proof of Proposition 1
$\underset{x\to a}{\lim}f(x,y_{0})$ $=\underset{y\to c}{\lim}g(x_{0},y)=0$
the existence of the solution follows easily from Poincaré–Miranda theorem. For the uniqueness of the solution, notice that by assumptions
it follows that the Jacobian of $h:\mathbb{R}^{2}\to\mathbb{R}^{2}$ defined by $h(x,y)=(f(x,y),g(x,y))$ is always positive:
$J_{h}(x,y)=\left|\begin{array}[]{cc}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}\\ \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}\end{array}\right|=\frac{\partial f}{\partial x}\cdot\frac{\partial g}{\partial y}-\frac{\partial f}{\partial y}\cdot\frac{\partial g}{\partial x}>0.$
Hence, $h$ is one-to-one, yielding that the solution is unique. This gives us the first part of Proposition 1.
For the second part, let us write the equations
@@
$\begin{array}[]{c}1=\frac{\partial f}{\partial C}=\frac{\partial f}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial f}{\partial x}\cdot\frac{\partial x}{\partial C}\\ 1=\frac{\partial g}{\partial C}=\frac{\partial g}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial g}{\partial x}\cdot\frac{\partial x}{\partial C}\end{array}\Rightarrow\left(\begin{array}[]{cc}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}\\ \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}\end{array}\right)\cdot\left(\begin{array}[]{c}\frac{\partial x}{\partial C}\\ \frac{\partial y}{\partial C}\end{array}\right)=\left(\begin{array}[]{c}1\\ 1\end{array}\right)$
@@ 6 Determining filters biases - demonstration
Then, we sampled the plane spanned by $L=$ bias_fo and $H=$ bias_hpf in steps of 10 units. For each sample we used an automatic feedback control algorithm that balanced the positive and negative event rate to be close to 50k events/sec and 100k events/sec in total. We remark that due to Proposition 1, the convergence of the algorithm to the desired background event rate was guaranteed. Conceptually, the algorithm performs a similar idea as the AGC feature in frame based cameras.
@@ 7 Conclusions
Solving the biases multi-variable problem is a crucial stage one needs to pass to exploit the potential of event camera for performing any task. Hence, developing systematic heuristics and approaches to tackle this problem is an important challenge of the research community. In this paper we presented a systematic heuristic to characterize the biases of an event camera for detection of small objects tasks in staring scenarios. We translated the biases behavior in this context into mathematical terms and showed that under suitable and reasonable assumptions, for any value of the filters biases, there exists exactly one pair of sensitivity biases that provide a predefined positive and negative event rate. Having this understanding in mind, we demonstrated, in a certain scenario, under appropriate constraints on the event rate, how the sensitivity biases change, and how the signal of an incandescent lamp powered by the electrical grid is being influenced as well. This demonstraion gave us a better understanding on what can be done with event camera when one wants to detect signals with similar behavior.